OpenStudy (anonymous):

all solutions to \(x^2+y=xy^2\) \(y^2+x=yx^2\) \(x\) , \(y\) are real numbers

OpenStudy (anonymous):

If we subtract them then it will work or not ??

OpenStudy (anonymous):

it will work

OpenStudy (anonymous):

\[(y-x)(-((x+y)+ 1) = xy(y-x)\]

OpenStudy (anonymous):

\[\implies 1 - x - y = xy\]

OpenStudy (anonymous):

S={(a;a)/ a in B}.....B={-1;0;(1+sqrt(5))/2;(1-sqrt(5))/2}

OpenStudy (anonymous):

@waterineyes u forgot the case \(y=x\)

OpenStudy (anonymous):

x^2 - y^2 - (x-y) = -xy(x-y) (x-y)(x+y -1+xy) = 0 => x+y-1+xy = 0 => y = (1-x)/(1+x) 1 + x =/= 0

OpenStudy (anonymous):

Yeah in that case we cannot divide both sides..

OpenStudy (anonymous):

x=1, y=0 y=1. x=0

OpenStudy (anonymous):

i can't get other points

OpenStudy (anonymous):

3 solutions from \(x=y\) and any from \(x+y+xy=1\)

OpenStudy (anonymous):

For x = y there are three solutions we will get @mukushla ?? Asking..

OpenStudy (anonymous):

yes...